# Post-detection inference for sequential changepoint localization

### v4

@@ C.1 A concrete example: Gaussian mean shift problem
C.2 Another concrete example: Poisson rate change problem
Here is $F_{0}$ is $\text{Pois}(\theta_{0})$ and the post-change model is $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\in\Theta_{1}\}$.
In this setting, we have the following example of a $1-\beta$ confidence sequence [2] for $\theta_{1}$
$\mathcal{S}^{\prime}_{n}=\{\theta:n\theta-\sum_{i=t}^{t+n-1}X_{i}\log\theta\leq\log\left(\frac{1}{\beta}\right)+\log I(c,c\theta)-\log I(n+c,\sum_{i=t}^{t+n-1}X_{i}+c\theta)\}\cap\Theta_{1},$ (S13)
where $c$ is any constant and $I(a,b)=\int_{-\infty}^{\infty}e^{-ae^{x}+bx}dx$.
We fix $t\in\{1,\cdots,\tau\}$. Let $\lambda_{t}=\max\{\theta_{0},\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$. We only need to generate $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Pois}(\lambda_{t})$ and $U_{n,1}^{j},\cdots,U_{n,X_{n}^{j}}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Uniform}(0,1)$, independently for each $n\in\mathbb{N},j=1,\cdots,B$. For $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$,
$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta_{0}/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ otherwise. }\end{cases}$ (S14)
Using the binomial thinning property of Poisson distribution, it is straightforward to verify that $Y_{n}^{t}(X_{n}^{j},\theta)\sim\text{Pois}(\theta_{0})$, for $n<t$ and $Y_{n}^{t}(X_{n}^{j},\theta)\sim\text{Pois}(\theta)$, for $\theta\in\mathcal{S}^{\prime}_{\tau-t+1},n\geq t$.
We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in Section 3, with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Let $\tau^{\theta}_{j}$ be the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$. Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark C.3) and they are finite. So, we can choose $L=\infty$.
If $t_{2}^{j}<t$, we have $V_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t,V_{t,\infty}^{j}$ can be simplified as below:
$V_{t,\infty}^{j}$ $=\max_{\begin{subarray}{c}i,t^{\prime}\in\mathbb{N}:1\leq i\leq t^{\prime},\\ \max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}\end{subarray}}\max_{\theta\in{S}^{j}_{t,t^{\prime}}}\frac{L_{i}\left(\frac{\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta)}{t^{\prime}-i+1};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}{L_{t}\left(\frac{\sum_{k=t}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta)}{t^{\prime}-t+1};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}.$
Here ${S}^{j}_{t,t^{\prime}}=\{\lambda_{t}U_{n,k}^{j}:k,n\in\mathbb{N},k\leq X_{n}^{j},t\leq n\leq t^{\prime},\inf\mathcal{S}^{\prime}_{\tau-t+1}\leq\lambda_{t}U_{n,i}^{j}\leq\sup\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\inf\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$ is the set consisting of one candidate element from is a finite set with at most $2+\sum_{n=t}^{t^{\prime}}X_{n}^{j}$ many elements. All the maximum are taken over a finite set; hence, $V_{t}^{j}$ can be computed easily.
Remark C.3.
For weighted CUSUM or SR detectors (defined in Remark 3.3.), finding $t_{1}^{j},t_{2}^{j}$ is often straightforward. For example, consider $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$. Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{(\tau-j+1)\times(\theta-\theta_{0})\}(\theta/\theta_{0})^{\sum_{i=j}^{\tau}X_{i}}$ increases as $X_{i}$’s increases and hence both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases. Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Similarly, for $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Also, for $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\neq\theta_{0}\}$, if we employ Gamma weights (it is a conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution.
@@ Appendix D Computationally and statistically efficient subroutines in Algorithm 3
Let $L_{t}^{n}(\theta_{0},\theta_{1})$ be the likelihood of data up to time $n\geq t$ under $H_{0,t}$. Then,
$L_{t}(\theta_{0},\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (S1)
We define $\hat{\theta}_{0,1:j}$ to be the maximum likelihood estimator (MLE) of the null $\theta_{0}\in\Theta_{0}$ based on data $X_{1},\cdots,X_{j}$. Similarly, we define $\hat{\theta}_{1,j:n}$ to be the MLE of the alternative $\theta_{1}\in\Theta_{1}$ based on data $X_{j},\cdots,X_{n}$. The point estimate of $T$ can be naturally extended to the following form:
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\hat{\theta}_{0,1:j-1}}(X_{i})}.$ (S2)
As a prototypical example, we consider the likelihood-ratio-based test statistic $M_{t}^{(n)}$, which is defined for $t,n\in\mathbb{N},t\leq n$ as below.
$M_{t}^{(n)}(X_{1},\cdots,X_{n}):=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{0,1:i-1},\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (S3)
Note that the test statistic $M_{t}$ for testing the null hypothesis $H_{0,t}$ reduces to
[…]
@@
Note that, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta,\theta^{\prime})\}_{n}$ across all $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution.
Let us now fix $t\in\{1,\cdots,\tau\}$. Computing the detection time $\tau^{j}_{\theta,\theta^{\prime}}$ for each $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$, and finding whether $\text{there exists }(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$ are both computationally infeasible, unless $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ are a finite sets.
However, it is often straightforward to compute $t_{1}^{j}=\displaystyle\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$ and $t_{2}^{j}=\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$. Also, we can often derive a suitable upper bound on the threshold, which does not require any approximation and is computationally and statistically efficient. E.g., we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$. Again, we bound $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})$ by $U_{t,L}^{j}$, where for $L<\infty$,
$U_{t,L}^{j}:=\begin{cases}-\infty,\penalty 10000\ \penalty 10000\ \quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t\leq t_{2}^{j}\leq L\\ \infty,\penalty 10000\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L.\end{cases}$ (S6)
And for $L=\infty$,
$U_{t,\infty}^{j}:=\begin{cases}-\infty,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \penalty 10000\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{otherwise.}\end{cases}$ (S7)
@@
Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 4.4.
Corollary D.1.
$\mathcal{C}^{\prime\prime}_{1}$ defined in (S8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.4.
In the next subsection, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with Poisson distribution in Appendix C of the Supplementary Material.
@@ D.1 A concrete example: Gaussian mean shift problem
We consider the most common changepoint model, where pre-change and post-change models are $\mathcal{P}_{i}=\{N(\theta,1):\theta\in\Theta_{i}\}$, for $i=0,1$. Then, we have the following $1-\gamma$ confidence interval for $\theta_{0}$ for each $t\geq 2$:
@@
$Y_{n}^{j}(\theta,\theta^{\prime})=\begin{cases}\theta+\epsilon_{n}^{j},\text{ if }n<t\\ \theta^{\prime}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (S10)
We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (S4), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Note that, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly even with $L=\infty$ for some specific detection algorithms; see Remark D.2. For $t_{2}^{j}<t,U_{t,\infty}^{j}=0$ and for $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
$\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j}+\theta_{0},\bar{\epsilon}_{t:t^{\prime}}^{j}+\theta_{1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}$
$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\displaystyle\max_{1\leq i\leq t^{\prime}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j},\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})}$
$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\displaystyle\max_{1\leq i\leq t^{\prime}}L_{i}(\bar{Y}_{1:i-1}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}),\bar{Y}_{i:t^{\prime}}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i});Y_{1}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}),\cdots,Y_{t^{\prime}}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j},\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})},$
where $\bar{\epsilon}_{i:t^{\prime}}^{j}=\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j},\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime})=\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{j}(\theta,\theta^{\prime}),$ and $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ is the one such pair that maximizes $L_{i}\left(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)$, or equivalently minimizes
$-2\log L_{i}\Big(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\dots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\Big)=$
$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i<t,\\[8.0pt] \text{constant},&\text{if }i=t,\\[8.0pt] \displaystyle\sum_{k=1}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}-\frac{i-t}{i-1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}+\frac{t-1}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i>t,\end{cases}$ (S11)
over $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. Here, constant means that it does not involve $\theta,\theta^{\prime}$. Note that $\delta_{t^{\prime}}^{i}:=\theta_{1,t^{\prime}}^{i}-\theta_{0,t^{\prime}}^{i}$ have simple closed-form expressions since (D.1) is a quadratic function of $\theta_{1}-\theta_{0}$. Then, we can take $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ to be any pair $(\theta,\delta_{t^{\prime}}^{i}+\theta)\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
Remark D.2.
@@
For weighted CUSUM or SR detectors, finding $t_{1}^{j},t_{2}^{j}$ is often straightforward. So, for some weight $W(\theta)$ such that $\int_{\theta\in\Theta_{1}}dW(\theta)=1$, the weighted CUSUM or SR detectors (with the Reverse Information Projection (RIPr) [17, 8, 16] being the representative of the pre-change class) are defined as
@@
$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (S12)
@@
$\tau_{\text{wsr-ripr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (S13)
@@
$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }$
$t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$

### v6

@@ Appendix D Computationally and statistically efficient subroutines in Algorithm 3
Let us now fix $t\in\{1,\cdots,\tau\}$ and consider the test statistic $M_{t}$ of the form in (9), with some $R^{(t)}_{n}=R^{(t)}_{n}(\{X_{n}\}_{n})$ and $S^{(t)}_{n}=S^{(t)}_{n}(\{X_{n}\}_{n})$.. Computing the detection time $\tau^{j}_{\theta,\theta^{\prime}}$ for each $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$, and finding whether $\text{there exists }(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$ are both computationally infeasible, unless $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ are a finite sets.
However, it is often straightforward to compute $t_{1}^{j}=\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$. Also, we can often derive a suitable upper bound on the threshold, which does not require any approximation and is computationally and statistically efficient. E.g., we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$. Again, we bound $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})$ by $U_{t,L}^{j}$, where
[…]
@@
Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 8.
Corollary 2.
$\mathcal{C}^{\prime\prime}_{1}$ defined in (S21) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 8.
In the next subsection, we present a concrete example of the Gaussian mean shift problem to clarify our method.
@@ D.1 A concrete example: Gaussian mean shift problem
We consider the most common changepoint model, where pre-change and post-change models are $\mathcal{P}_{i}=\{N(\theta,1):\theta\in\Theta_{i}\}$, for $i=0,1$, with $\Theta_{0}=\{\theta:\theta\leq\theta_{0}\}$ and $\Theta_{1}=\{\theta:\theta\geq\theta_{1}\}$, for some $\theta_{1}>\theta_{0}$. Then, we have the following $1-\gamma$ confidence interval for $\theta_{0}$ for each $t\geq 2$:
@@
$Y_{n}^{t,j}(\theta,\theta^{\prime})=\begin{cases}\theta+\epsilon_{n}^{j},\text{ if }n<t\\ \theta^{\prime}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (S23)
We consider the test statistic $M_{t}$, as defined in (9), with $R_{n}^{(t)}=\int_{\theta\in\Theta_{0}}\prod_{i=t}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{1}}(X_{i})}w^{\prime}(\theta)d\theta,$ and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)d\theta$, with $w,w^{\prime}\geq 0$ being some weight functions such that $\int_{\theta\in\Theta_{1}}w(\theta)d\theta=1$ and $\int_{\theta\in\Theta_{0}}w^{\prime}(\theta)d\theta=1$. Note that $t_{1}^{j}$ is easy to compute for some CUSUM/SR type detection algorithms (see Remark 6) and they are finite. So, we can have $L=\infty$. Note that, here the term $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}R^{(t)}_{n}\left(\{Y_{n}^{t,j}(\theta,\theta^{\prime})\}_{n}\right)$ simplifies to
$\underset{\theta\in\Theta_{0}}{\int}\exp\left((\theta-\theta_{1})\sum_{i=t}^{n}(\epsilon_{n}^{j}+\theta^{\prime}_{t,\min})+\frac{n-t+1}{2}(\theta_{1}^{2}-\theta^{2})\right)w^{\prime}(\theta)d\theta,$
and the term $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}S^{(t)}_{n}\left(\{Y_{n}^{t,j}(\theta,\theta^{\prime})\}_{n}\right)$ simplifies to
$\underset{\theta\in\Theta_{1}}{\int}\exp\left((\theta-\theta_{0})\sum_{i=n}^{t-1}(\epsilon_{n}^{j}+\theta^{\prime}_{t,\min})+\frac{t-n}{2}(\theta_{0}^{2}-\theta^{2})\right)w(\theta)d\theta,$
where $\theta^{\prime}_{t,\min}=\inf\mathcal{S}^{\prime}_{\tau-t+1}$. And hence, $U_{t,\infty}^{j}$ in (S20) can be computed easily.
Remark 6.
@@
For weighted CUSUM or SR detectors, finding $t_{1}^{j}$ is often straightforward. So, for some weight $w(\theta)$ such that $\int_{\theta\in\Theta_{1}}w(\theta)=1$, the weighted CUSUM or SR detectors (with the Reverse Information Projection (RIPr) [larsson2024numeraire, grunwald2024safe, lardy2023universal] being the representative of the pre-change class) are defined as
@@
$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}w(\theta)\geq A\right\},$ (S24)
@@
$\tau_{\text{wsr-ripr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}w(\theta)\geq A\right\},$ (S25)
@@
$t_{1}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$